OpenStudy (anonymous):

Which polynomial is prime? a.4x^2-17x+15 b.8x^2+7x-16 c.4x^2+13x d.x^2-4

OpenStudy (anonymous):

Over the integers, the rationals?

OpenStudy (anonymous):

B over the rational

OpenStudy (anonymous):

you need to put them in the form ax^2 + bx + c = 0 if the descriminate (b^2-4ac) is a perfect square, then it's not prime..

OpenStudy (anonymous):

so wouldn't it be a?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

a can be factored the descriminant is 49

OpenStudy (anonymous):

\[ 4 x^2-17 x+15=(x-3) (4 x-5) \]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

c, cause its not factor able right

OpenStudy (anonymous):

am I right?

OpenStudy (anonymous):

I meant b! not c

OpenStudy (anonymous):

b is the answer